d. Calculate the following different elasticities:

i. A price increase from P2 to P10 causes quantity demanded to change from 80 units to 30 units. Calculate and interpret price elasticity of demand. (3 marks)

ii. Income increase by 10% results in quantity demanded increases by 5%. Calculate and interpret income elasticity of demand. (2 marks)

iii. Quantity of good B increases by 50% because of an increase in price of good A by 40%. Calculate and interpret cross elasticity of demand. (2 marks)

iv. How would a firm manipulate the prices of goods in di-iii in order to maximize revenue and profits?
Answer #1

i. Price elasticity of demand= (30-80)/(30+80)÷(10-2)/(10+2)= (-50/110)÷(8/12)= -0.68

Simce the absolute value is 0.68 which is less than 1, the price elasticity of demand is inelastic. It means consumers are unresponsive to the price changes.

ii. Income elasticity of demand= 5%/10%= 0.5

Since the income elasticity of demand is positive it means the income is elastic and consumer demand more of the good as the income increases.

iii. Cross elasticity of demand= 50%/40%= 1.25

Since the cross elasticity of demand ia positive, it means the good A and good B are substitute goods.

iv. di. When Price elasticity of demand is inelastic, increasing the price will lead to increase in total revenue.

dii. Even though income elasticity is not related to price but positive income elasticity means the good is a normal good. It can mean rising price leads to rising demand and revenues.

diii. Cross elasticity of demand is positive and goods are substitutes. When price of good A rises, demand for good B rises. So, When price of good A rises, the profit and revenue for good B also rises.